OpenStudy (anonymous):

How do I do the step by step factoring of this expression? 6(x + 5)^2 – 5(x + 5) + 1

OpenStudy (anonymous):

think of it as \(6u^2-5u+1\), factor that one, then replace \(u\) by \(x+5\)

OpenStudy (anonymous):

i think it factors as \((2u-1)(3u-1)\) but i am lousy at factoring so check it then get \((2(x+5)-1)(3(x+5)-1)\) and clean up inside the parentheses

OpenStudy (anonymous):

My teacher said the final answer was (2x + 9)(3x + 14) but Im having trouble showing my work to factor it out to get that answer.

OpenStudy (anonymous):

\[2(x+5)-1=2x+10-1=2x+9\] for the first parentheses \[3(x+5)-1=3x+15-1=3x+14\] for the second parentheses

OpenStudy (anonymous):

Ok, I think I see what I did wrong. Thanks for your help.